# VideoRoPE: What Makes for Good Video Rotary Position Embedding?

### v1

@@ Appendix F Supplementary Explanation on Frequency Allocation
(a) For M-RoPE (Wang et al., 2024b), temporal dependency is modeled using the first 16 rotary angles, which exhibit higher frequency and greater oscillation. Taking the last 3 rotary angles as an example, the position embedding for temporal modeling undergoes significant distortion due to periodic oscillations (Men et al., 2024), as these dimensions have shorter monotonous intervals. Lower dimensions have even shorter intervals. Notably, because the oscillation is periodic, two distant positions can have nearly identical position embeddings, resembling a hash collision, as shown by the red planes. This phenomenon is why distractors can easily mislead the model.

### v2

@@ Appendix F Supplementary Explanation on Frequency Allocation
(a) For M-RoPE (Wang et al., 2024a), temporal dependency is modeled using the first 16 rotary angles, which exhibit higher frequency and greater oscillation. Taking the last 3 rotary angles as an example, the position embedding for temporal modeling undergoes significant distortion due to periodic oscillations (Men et al., 2024), as these dimensions have shorter monotonous intervals. Lower dimensions have even shorter intervals. Notably, because the oscillation is periodic, two distant positions can have nearly identical position embeddings, resembling a hash collision, as shown by the red planes. This phenomenon is why distractors can easily mislead the model.